OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ?

OpenStudy (anonymous):

x = -7 or x = 5

OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ?

OpenStudy (anonymous):

either x=-7 or x=5

hero (hero):

x^2 + 2x -35 = 0 x^2+7x - 5x - 35 = 0 x(x+7) -5(x+7) = 0 (x+7)(x-5) =0 x = -7 x = 5

OpenStudy (anonymous):

solving it by middle term breaking

OpenStudy (anonymous):

it's a binomial that you have to factor. x^2 +2x-35=0 (x+7)(x-5) then you set them equal to 0. so x=-7 or x=5

OpenStudy (anonymous):

(1)x*x + (2)x + (-35) = 0 a = 1 b = 2 c = -35 (-b +/- d)/2a d*d = b*b - 4ac x = -5, 7

OpenStudy (anonymous):

hero,s method is rite

OpenStudy (anonymous):

-7 and 5 u r wrong fibonacci girl

OpenStudy (anonymous):

check again

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

yes lol :)

OpenStudy (anonymous):

my ans is rite